A 4.0kg object has a velocity of 3.0i m/s at one instant. Eight seconds later, its velocity is (8.0i+10.0j) m/s . Assuming the object was subject to a constant net force, find (a) the components of the force and (b) its magnitude.
1
Expert's answer
2021-08-10T12:16:04-0400

The given information is

  • The initial velocity is "V_{i}=3.0\;i\;\text{m}/\text{s}"
  • The final velocity is "V_{f}=(8.0\;i+10.0\;j)\;\text{m}/\text{s}"
  • The time it takes to change velocity is "t=8\;\text{s}"
  • The mass of the object is "m=4.0\;\text{Kg}"

Part A


The acceleration is given by.

"a=\dfrac{V_{f}-V_{i}}{t}"

where,

"V_{i}" Is the initial velocity.

"V_{f}" Is the final velocity.

"t" Is the time.


Evaluating numerically.

"a=\dfrac{V_{f}-V_{i}}{t}\\
.\\
a=\dfrac{(8.0\;i\;+10.0\;j)\;\text{m}/\text{s}-3.0\;i\;\text{m}/\text{s}}{8.0\;\text{s}}\\
.\\
a=\dfrac{8.0\;i\;\text{m}/\text{s}+10.0\;j\;\text{m}/\text{s}-3.0\;i\;\text{m}/\text{s}}{8.0\;\text{s}}\\
.\\
a=\dfrac{5.0\;i\;\text{m}/\text{s}+10.0\;j\;\text{m}/\text{s}}{8.0\;\text{s}}\\
.\\
a=\dfrac{(5.0\;i+10.0\;j)\;\text{m}/\text{s}}{8.0\;\text{s}}\\
.\\
a=(0.625\;i+1.25\;j)\;\text{m}/\text{s}^{2}\\"


The force on the object is given by.

"F=m\;a\\"


where.

"m" Is the mass.

"a" is the acceleration.


Evaluating numerically.

"F=m\;a\\
F=4.0\;\text{Kg}\times (0.625\;i+1.25\;j)\;\text{m}/\text{s}^{2}\\
F=(2.5\;i+5\;j)\;\text{N}"


Answer (a)

The force on the object is 

The horizontal component. "\displaystyle \color{red}{\boxed{F_{x}=2.5\;\text{N}}}"

The vertical component. "\displaystyle \color{red}{\boxed{F_{y}=5\;\text{N}}}"


Part B


The magnitude of the force is given by.

"F=\sqrt{ F_{x}^{2}+F_{y}^{2}}"


Evaluating numerically. 

"F=\sqrt{ F_{x}^{2}+F_{y}^{2}}\\
F=\sqrt{ (2.5\;\text{N})^{2}+(5\;\text{N})^{2}}\\
F=\sqrt{ 31.25\;\text{N}^{2} }\\
F=5.6\;\text{N}"


Answer B

The magnitude of the force on the object is "\displaystyle \color{red}{\boxed{F=5.6\;\text{N}}}"